OpenStudy (davidusa):

Is there a trick to quickly solving magic squares? MEDAL AND FAN FOR BEST HELPER!

OpenStudy (anonymous):

magic squares?

OpenStudy (anonymous):

Magic is fake.

thomaster (thomaster):

By the way, the trick is: \(\LARGE M=\dfrac{n(n^2+1)}{2}\)

thomaster (thomaster):

So if you have a magic square of 3 x 3 you'd have \(\LARGE \dfrac{3(3^2+1)}{2}=15\)

thomaster (thomaster):